Two charged objects separated by 21 cm have a repulsive force of 0.02 N. If the distance in charged to 7 cm, what will be be the resulting repulsive force?
1
Expert's answer
2011-01-31T12:25:59-0500
The repulsive force is
F1 = kq1q2/R12

kq1q2 = F1R12
Then
F2 =k q1q2/R22 = F1R12/R22 = 0.02*212/72 = 0.18 N.